4(1.5c + 6) - 2c = -9
Answer by anjali_yamini(27)


4(1.5c + 6) -2c = -9
6.0c + 24 -2c = - 9
6c -2c +24 = -9
4c = -9 -24
4c = - 33
c = -33 / 4
c = - 3.75